# Polyhedral Complex Derivation from Piecewise Trilinear Networks

### v1

@@ Appendix C Experimental Results
Marching Cubes   16     303 $\pm$       1 13643 $\pm$   55 4.21 0.00 $\pm$ 0.00
  24     704 $\pm$     14   8576 $\pm$ 173 4.39 0.00 $\pm$ 0.00
  32   1338 $\pm$     26   5842 $\pm$ 198 4.68 0.00 $\pm$ 0.00
  40   2105 $\pm$     41   5134 $\pm$ 156 4.25 0.01 $\pm$ 0.00
  48   3065 $\pm$     39   4400 $\pm$   81 4.11 0.04 $\pm$ 0.06
  56   4202 $\pm$     62   3896 $\pm$ 136 3.96 0.01 $\pm$ 0.00
  64   5570 $\pm$     78   3252 $\pm$ 733 4.12 0.02 $\pm$ 0.00
128 22612 $\pm$   378   3053 $\pm$ 102 2.18 0.10 $\pm$ 0.00
196 51343 $\pm$   826   1471 $\pm$ 745 3.00 0.34 $\pm$ 0.01
224 70007 $\pm$ 1128   1614 $\pm$ 473 2.34 0.62 $\pm$ 0.02
Marching ${}^{\star}$ 256 91641 $\pm$ 1521         0 $\pm$     0 - 0.89 $\pm$ 0.03
Tropical (ours) -   4491 $\pm$     92   2593 $\pm$ 484 5.75 5.66 $\pm$ 1.40
[…]
Method Sample Vertices $\downarrow$ AD (${}^{\circ}$) $\downarrow$
Marching Cubes   16     303 $\pm$       1 16.0 $\pm$ 12.5
  24     704 $\pm$     14 13.2 $\pm$ 10.6
  32   1339 $\pm$     27   9.8 $\pm$   8.7
  40   2107 $\pm$     41   8.2 $\pm$   7.3
  48   3067 $\pm$     39   6.8 $\pm$   6.7
  56   4205 $\pm$     63   5.6 $\pm$   6.2
  64   5569 $\pm$     77   4.9 $\pm$   5.2
128 22615 $\pm$   380   2.9 $\pm$   4.0
196 51345 $\pm$   826   1.9 $\pm$   2.9
224 70015 $\pm$ 1131   1.8 $\pm$   2.8
Marching ${}^{\star}$ 256 91643 $\pm$ 1521 -
Tropical (ours) -   4490 $\pm$     91   3.8 $\pm$   4.5
Appendix D Visualizations
@@
In Figure 7, we showcase the analytical generation of a normal map from piecewise trilinear networks, utilizing HashGrid (Müller et al., 2022) and ReLU neural networks. Our approach dynamically assigns vertices based on the learned decision boundary. Notably, the configuration of hash grids influences vertex selection, and each grid can be subdivided into multiple polyhedra for accurate representation of smooth curves.
In Figure 8, we present a comparative visualization of normal maps for marching cubes, demonstrating variations in the number of samplings alongside our method, with the respective vertex counts indicated in parentheses. This analysis highlights the trade-off between detail preservation and computational efficiency, showcasing the optimality of our analytically extracted mesh from decision boundary apex points.
@@
Figure 7: The normal map is analytically generated from piecewise trilinear networks, comprising both HashGrid (Müller et al., 2022) and ReLU neural networks, that have been trained to learn a signed distance function with the eikonal loss for the Stanford bunny (Curless & Levoy, 1996). Our approach dynamically assigns vertices based on the learned decision boundary. One notable observation is that the configuration of hash grids influences the selection of vertices. Additionally, each grid can be subdivided into multiple polyhedra to accurately represent smooth curves.

### v2

@@ Appendix F Experimental results
Marching Cubes   32   1283 $\pm$     89   4106 $\pm$ 202 19.0 0.00 $\pm$ 0.00
  48   2967 $\pm$   177   2178 $\pm$   94 15.5 0.01 $\pm$ 0.06
  56   4057 $\pm$   245   1539 $\pm$ 100 16.0 0.01 $\pm$ 0.00
  64   5367 $\pm$   342   1371 $\pm$   93 13.6 0.02 $\pm$ 0.00
128 21825 $\pm$ 1336     393 $\pm$   19 11.6 0.10 $\pm$ 0.00
192 49569 $\pm$ 3043     141 $\pm$    4 14.3 0.44 $\pm$ 0.10
224 67601 $\pm$ 4146     114 $\pm$    6 12.9 0.60 $\pm$ 0.03
Marching Cubes 256 88492 $\pm$ 5404            - - 0.88 $\pm$ 0.05
Ours -   4341 $\pm$   253     900 $\pm$ 171 25.6 0.97 $\pm$ 0.21
Table 5: Angular distance (AD) for the Stanford bunny. Similarly to the chamfer distance, the angular distance () is also evaluated using the marching cubes with 2563 grid samples ($\star$) as the reference ground truth for all measurements. The standard deviations are provided after the $\pm$ notation, based on results obtained from three random seeds, which is used in training.
Method Sample Vertices $\downarrow$ AD () $\downarrow$
Marching Cubes   32   1283 $\pm$     89   6.97 $\pm$ 0.45
  48   2967 $\pm$   177   5.07 $\pm$ 0.15
  56   4057 $\pm$   245   4.23 $\pm$ 0.25
  64   5367 $\pm$   342   4.10 $\pm$ 0.17
128 21825 $\pm$ 1336   2.47 $\pm$ 0.12
192 49569 $\pm$ 3043   1.63 $\pm$ 0.06
224 67601 $\pm$ 4146   1.53 $\pm$ 0.06
Marching Cubes 256 88492 $\pm$   253 -
Ours -   4490 $\pm$     91   3.57 $\pm$ 0.38
Table 6: Chamfer distance (CD) and chamfer efficiency (CE) for the Part A of the Stanford 3D Scanning repository [30] using the Large model having 4x-resolution of HashGrid compared to the default setting (ref. Section 6). The chamfer distance ($\times$1e-6) is evaluated using the marching cubes (MC) with 5123 grid samples as the reference ground truth for all measurements. The results are averaged over three random seeds.
Bunny Dragon Happy Buddha
MC # $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$
128   38395 632 4.12   25986 759 5.07 18621 1110 4.84
192   87245 299 3.83   58855 335 5.08 42793   533 4.38
224 119000 221 3.80   80470 268 4.64 58357   395 4.34
256 155484 182 3.53 105235 226 4.21 76183   322 4.07
Ours 135691 151 4.87   91147 157 6.99 67562   229 6.46
Table 7: Chamfer distance (CD) and chamfer efficiency (CE) for the Part B of the Stanford 3D Scanning dataset [30] using the Large model as the default setting (ref. Section 6).
Armadillo Drill Bit Lucy
MC # $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$
128   25309 887 4.45   3590 1342 20.8 12224 1304 6.28
192   57751 399 4.34   8092   625 19.8 27863   687 5.23
224   78703 315 4.04 11000   588 15.5 38337   524 4.97
256 103085 267 3.63 14339   496 14.1 49838   435 4.61
Ours   90567 186 5.94 10275   342 28.5 43631   358 6.40
Table 8: Chamfer distance (CD) and chamfer efficiency (CE) for the Stanford dragon using the Large model having 4x-resolution of HashGrid compared to the default setting (ref. Section 6) are evaluated using the marching cubes with 5123 grid samples ($\star$) as the reference ground truth.
Marching Cubes 128   25986 $\pm$   121     759 $\pm$   64   5.07 0.29 $\pm$ 0.01
192   58855 $\pm$   181     335 $\pm$   21   5.08 0.60 $\pm$ 0.06
256 105235 $\pm$   452     226 $\pm$   17   4.21 1.16 $\pm$ 0.04
Marching Cubes 512 424311 $\pm$ 1262            - - 9.19 $\pm$ 0.39
Ours -   91147 $\pm$   203     157 $\pm$   18   6.99 53.5 $\pm$ 16.3
Appendix G Visualizations
@@
In Figure 5, we showcase the analytical generation of a normal map from piecewise trilinear networks, utilizing HashGrid [12] and ReLU neural networks. Our approach dynamically assigns vertices based on the learned decision boundary. Notably, the configuration of hash grids influences vertex selection, and each grid can be subdivided into multiple polyhedra for accurate representation of smooth curves.
In Figure 6, we present comparative visualizations of normal maps for marching cubes, demonstrating variations in the number of samplings alongside our method, with the respective vertex counts indicated in parentheses. This analysis highlights the trade-off between detail preservation and computational efficiency, showcasing the optimality of our analytically extracted mesh from decision boundary apex points.
@@
In Figure 8, the gallery of the six Stanford 3D Scanning meshes from the Large models is shown.
@@
Figure 5: The normal map is analytically generated from piecewise trilinear networks, comprising both HashGrid [12] and ReLU neural networks, that have been trained to learn an SDF with the eikonal loss for the Stanford bunny [30]. Our approach dynamically assigns vertices based on the learned decision boundary. One notable observation is that the hash grids influence the selection of vertices. Additionally, each grid can be subdivided into multiple polyhedra to accurately represent smooth curves. The Small (left) and Large (right) models (ref. Section 6) are used to learn the SDF.
@@
Figure 8: The extract mesh visualizations of the six from the Stanford 3D Scanning repository using the Large models. From left, Bunny, Dragon, Happy Buddha, Armadillo, Dril Bit, and Lucy.